OpenStudy (anonymous):

express z=(sqrt3-i)^7/(sqrt2+isqrt2)^6 in trig form z=p(cos theta + i sin theta)

OpenStudy (anonymous):

I not sure where to start should I convert the sqrt3-i into polar form?

zepdrix (zepdrix):

Convert the numerator and denominator into polar separately.

zepdrix (zepdrix):

\[\Large\rm \sqrt3-1i\]\[\Large\rm r=\sqrt{(\sqrt3)^2+(1)^2}=\sqrt{10}\]\[\Large\rm \tan \theta=\frac{-1}{\sqrt3}\implies\quad \theta=\frac{5\pi}{6}\]So in the numerator we have:\[\Large\rm \left(\sqrt{10}\left(\cos\frac{5\pi}{6}+i \sin\frac{5\pi}{6}\right)\right)^7\]From there we would apply De' Moivre's Theorem to deal with the exponent. I skipped over a few steps in there, any confusion on finding the radial distance or the angle?

zepdrix (zepdrix):

Woops, that r should be... 3+1, sqrt(4)=2.

OpenStudy (anonymous):

that's what i got for the numerator =2, so , the denominator look like it's already in polar form

zepdrix (zepdrix):

It does? 0_o hmm

zepdrix (zepdrix):

So lemme wrap up the numerator a sec:\[\Large\rm =\left(2\left[\cos\frac{5\pi}{6}+i \sin\frac{5\pi}{6}\right]\right)^7\] \[\Large\rm =2^7\left[\cos\frac{35\pi}{6}+i \sin\frac{35\pi}{6}\right]\]So this is our numerator in polar form ^ We would like to do similar to the denominator

zepdrix (zepdrix):

Can you apply those same steps to the denominator? We need a radius, and an angle.

zepdrix (zepdrix):

..? Confused? What's going on in the brain there? :o

OpenStudy (anonymous):

ohh no sorry computer was freezing lol , so denominator i got 2^6(cos 45deg(pi/4)+i sin 45deg)^6

OpenStudy (anonymous):

then what would be the next step once both are in polar form

zepdrix (zepdrix):

Good good good. So apply De' Moivre's Theorem gets us to this part of the problem:\[\Large\rm =\frac{2^7\left(\cos\frac{35\pi}{6}+i \sin\frac{35\pi}{6}\right)}{2^6\left(\cos\frac{6\pi}{4}+i \sin\frac{6\pi}{4}\right)}\]

zepdrix (zepdrix):

I can't remember what this rule is called.. but you're allowed to this:\[\Large\rm \frac{r_1(\cos a+i \sin a)}{r_2(\cos b+i \sin b)}=\frac{r_1}{r_2}\left[\cos(a-b)+i \sin(a-b)\right]\]We subtract the angles, and divide the radial lengths.

zepdrix (zepdrix):

It follows from the fact that these polar expressions can be written as exponentials. And the exponential law that we're familiar with (dividing exponentials), carries over to complex numbers.

OpenStudy (anonymous):

soo this formula can be used, i thought the r1/r2 formula was for two completly separate equations

zepdrix (zepdrix):

If we converted to exponentials, it might make more sense why this rule can be applied:\[\Large\rm \frac{2^7\left(\cos\frac{35\pi}{6}+i \sin\frac{35\pi}{6}\right)}{2^6\left(\cos\frac{6\pi}{4}+i \sin\frac{6\pi}{4}\right)}=\frac{2^7e^{(35\pi /6)i}}{2^6e^{(6\pi/4)i}}\]And of course, our rules of exponents would tell us that this is:\[\Large\rm \frac{2^7}{2^6}e^{\left(\frac{35\pi}{6}-\frac{6\pi}{4}\right)i}\]

OpenStudy (anonymous):

ohh ok makes sense when you use e

zepdrix (zepdrix):

Shouldn't be too bad from there, right? :) The 2's divide nicely. Just need to find a common denominator so you can subtract.

OpenStudy (anonymous):

yes thanks a lot

zepdrix (zepdrix):

np